OpenStudy (anonymous):

how do you complete the square for 4y^2-4y-35? with out factoring

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Completing_the_square check out the non-monic case but that seems like an awkward way to solve this particular equation

OpenStudy (anonymous):

yea i can do it when i factor out the 4 but then i need to put it in the non factored form so that i put it in the derivates of tan-1(x)

OpenStudy (anonymous):

(2y-1)^2 - 36 ?

OpenStudy (anonymous):

how did you get that? can you pls show me the steps?

OpenStudy (anonymous):

\[4y^2-4y-35=0\] \[( ? y - ? )^2 = 35\] \[(2y-1)^2-1=35\] \[(2y-1)^2-36=0\]

OpenStudy (anonymous):

thanks :)